Circular No. 6237

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ab IN NGC 7663 C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of an apparent supernova (V = 17.7) on a technical pan film obtained on Sept. 22.97 UT with the 0.9-m OCA Schmidt telescope; the supernova is located at α = 23^{h}26^{m}44^{s}.46, δ = −4^{o}57′45′′.9 (equinox 2000.0), which is 11′′.3 west and 14′′.1 north of the spiral galaxy’s center. SN 1995ab is not visible at this location on the Palomar Sky Survey prints. A nearby star (V ∼ 13.5) has position end figures 47^{s}.77, 45′′.5.

A. V. Filippenko and D. C. Leonard, University of California at Berkeley, report that CCD spectra (range 310–1000 nm, resolution 0.7–1.4 nm) obtained on Sept. 26 UT with the 3-m Shane reflector at Lick Observatory show that this supernova is of type II. The prominent, broad hydrogen Balmer lines have P-Cyg profiles. The narrow emission lines of a superposed H II region yield a redshift of 0.019.

SUPERNOVA 1995ac IN ANONYMOUS GALAXY Pollas also reports his discovery, on an OCA technical pan film taken on

Sept. 22.91 UT, of a supernova (V ∼ 18.5) located 2′′.5 west and 4′′.2 south of the center of an edge-on spiral galaxy (mag ∼ 16.5) at α = 22^{h}45^{m}41^{s}, δ = −8^{o}45′.2 (equinox 2000.0). Nothing is visible at this location on the PSS.

Filippenko and Leonard report that CCD spectra obtained as above (range 350–1030 nm) show that SN 1995ac closely resembles the peculiar type-Ia supernova 1991T (e.g., Filippenko et al. 1992, Ap.J. 384, L15), within a week prior to maximum brightness. The deep Si II absorption trough normally seen at 615 nm is absent; instead, there are blends (probably of Fe III) near rest wavelengths of 420 and 490 nm. The likely redshift of the parent galaxy, 0.05, was determined from the weak Hα emission line in the galaxy disk.

SUPERNOVA 1995Y IN NGC 410 P. Garnavich, A. Riess, P. Challis, and R. Kirshner, Center for Astro-

physics, report: “Spectra obtained by J. Huchra with the 1.5-m Tillinghast telescope on Sept. 22 show that this is a type-Ia supernova, 5–7 weeks after maximum. This is significantly younger than the age reported on IAUC 6221 and implies the discovery by Mueller was made soon after maximum. Photometry from CCD images obtained by B. McLean with the F.L. Whipple Observatory 1.2-m telescope yields: Sept. 14.4 UT, V = 16.96 ± 0.08; 19.4, V = 17.21 ± 0.05, B − V = +1.22 ± 0.07.”

1995 September 26 Daniel W. E. Green